OpenStudy (asnaseer):

Puzzle: Who can spot what is wrong here? Theorem: 4 = 5 Proof: -20 = -20 16 - 36 = 25 - 45 4^2 - (9x4) = 5^2 - (9x5) 4^2 - (9x4) + 81/4 = 5^2 - (9x5) + 81/4 (4 - 9/2)^2 = (5 - 9/2)^2 4 - 9/2 = 5 - 9/2 4 = 5

OpenStudy (lgbasallote):

how did you get -20=-20 from 4=5

OpenStudy (anonymous):

It's a puzzle.

OpenStudy (asnaseer):

The Theorem states 4 = 5 It is then followed by the proof

OpenStudy (lgbasallote):

oh that

OpenStudy (anonymous):

the third to last step

OpenStudy (anonymous):

sqrt of neg?

OpenStudy (asnaseer):

and what exactly is wrong with that step?

OpenStudy (anonymous):

The step is squaring a negative 1/2 and the other is squaring 1/2

OpenStudy (asnaseer):

which is fine, since (-1/2)^2 = (1/2)^2

OpenStudy (anonymous):

Yes but then the square root of that after is where the problem is

OpenStudy (asnaseer):

correct :)

OpenStudy (anonymous):

you're choosing the positive side

OpenStudy (anonymous):

not the same negative part you're squaring

OpenStudy (cwrw238):

yes - well done frees

OpenStudy (asnaseer):

it is essentially the same as saying: (-1)^2 = (1)^2 therefore -1 = 1

OpenStudy (asnaseer):

well spotted!

OpenStudy (anonymous):

\[\pm[4-\frac92]=\pm[5-\frac92]\]